The covariance of X and Y is 1. E(X) =3, E(Y) = 2. What is E(3XY)?

21

Not enough information

18

7
Answer #1

By the formula for covariance:

Cov(XY) = E(XY) - E(X)E(Y)

So,

E(XY) = Cov(XY) + E(X)E(Y) = 1 + 3*2 = 7

Hence,

E(3XY) = 3*E(XY) = 3*7 = 21